40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 13 10:54:55 2018, aaron, Update, VAC, Vent 80 vent80wtiptilts.png
    Reply  Fri Jul 13 16:26:52 2018, Steve, Update, VAC, Vent 80 is completing... @atm.pngvent80_7h.pngventregN2&Air_c.jpg
Message ID: 14064     Entry time: Fri Jul 13 10:54:55 2018     Reply to this: 14066
Author: aaron 
Type: Update 
Category: VAC 
Subject: Vent 80 

[aaron, steve]

Steve gave me a venting tutorial. I'll record this in probably a bit more detail than is strictly necessary, so I can keep track of some of the minor details for future reference.

Here is Steve's checklist:

  • Check that all jam nuts are tightened
  • all viewports are closed
  • op levs are off
  • take a picture of the MEDM screens
  • Check particle counts
  • Check that the cranes work & wiped
  • Check that HV is off

Gautam already did the pre-vent checks, and Steve took a screenshot of the IFO alignment, IMC alignment, master op lev screen, suspension condition, and shutter status to get a reference point. We later added the TT_CONTROL screen. Steve turned off all op levs.

We then went inside to do the mechanical checks

  • N2 cylinders in the 40m antechamber are all full enough (have ~700psi/day of nitrogen)
  • We manually record the particle count
    • this should be <10,000 on the 0.5um particles to be low enough to vent, otherwise we will contaminate the system
    • note: need to multiply the reading on the particle counter by 10 to get the true count
    • the temperature inside the PSL enclosure should be 23-24C +/- 3 degrees
    • We recorded the particle counts at ~830 and ~930, and the 0.5um count was up to ~3000
  • We put a beam stop in front of the laser at the PSL table
  • Checked that all HV supplies are either off or supplying something in air
    • we noticed four HV supplies on 1X1 that were on. Two were accounted for on the PSL table (FSS), and the other two were for C1IOO_ASC but ran along the upper cable rack. We got ahold of Gautam (sorry!) and he told us these go to the TT driver on OMC_SOUTH, where we verified the HV cables are disconnected. We took this to mean these HV supplies are not powering anything, and proceeded without turning these HV off.
    • There are HV supplies which were all either off or supplying something in-air at: 1Y4, 1Y2, OMC N rack, 1X9 (green steering HV)
  • Checked that the crane works--both move up and down
    • vertex crane switch is on the wall at the inner corner of the IFO
    • y arm crane switch is on the N wall at the Y end
    • turn off the cranes at the control strip after verifying they work
  • While walking around checking HV, we checked that the jam nuts and viewports are all closed
    • we replaced one viewport at the x arm that was open for a camera

After completing these checks, we grabbed a nitrogen cylinder and hooked it up to the VV1 filter. Steve gave me a rundown of how the vacuum system works. For my own memory, the oil pumps which provide the first level of roughing backstream below 500mtorr, so we typically turn on the turbo pumps (TP) below that level... just in case there is a calibrated leak to keep the pressure above 350mtorr at the oil pumps. TP2 has broken, so during this vent we'll install a manual valve so we can narrow the aperture that TP1 sees at V1 so we can hand off to the turbo at 500mtorr without overwhelming it. When the turbos have the pressure low enough, we open the mag lev pump. Close V1 if things screw up to protect the IFO. This 6" id manual gatevalve will allow us throttle the load on the small turbo while the maglev is taking over the pumping  The missmatch in pumping speed is 390/70 l/s [ maglev/varian D70 ]  We need to close down the conductive intake of the TP1 with manual gate valve so the 6x smaller turbo does not get overloaded...

We checked CC1, which read 7.2utorr.

Open the medm c0/ce/VacControl_BAK.adl to control the valves.

Steve tells me we are starting from vacuum normal state, but that some things are broken so it doesn't exactly match the state as described. In particular, VA6 is 'moving' because it has been disconnected and permanently closed to avoid pumping on the annulus. During this v ent, we will also keep pumping on the RGA since it is a short vent; steve logged the RGA yesterday.

We began the vent by following the vacuum normal to chamber open procedure.

  1. VM1 closed
  2. We didn't open VM3, because we want to keep the RGA on
  3. Closed V1
  4. Connect the N2 to the VV1 filter
    1. first puged the line with nitrogen
    2. We confirmed visually that V1 is closed
  5. We opened VM2 to pump on the RGA with the mag lev pump.
    1. This is a nonstandard step because we are keeping the RGA pumped down.
    2. The current on TP3 is ~0.19A, which is a normal, low load on the pump
  6. VV1 opened to begin the vent at ~10:30am
    1. use crescent wrench to open, torque wrench wheel to close
    2. Keep the pressure regulator below 10 psi for the vent. We started the vent with about 2psi, then increased to 8psi after confirming that the SUS sensors looked OK.
  7. We checked the pressure plot and ITMX/ETMX motion to make sure we weren't venting too quickly or kicking the optics
    1. Should look at eg C1:SUS-ITMX_SENSOR_LL, as well as C1:Vac-P1_pressure
  8. Once the pressure reaches 25torr, we switched over to dry air
    1. wipe off the outside dolly wheels with a wet rag, and exit through the x-arm door to get the air. Sweep off the area outside the door, and wipe off new air containers with the rag.
    2. Bring the cylinder inside, get the regulator ready/purged, and swap relatively quickly.
    3. We increased the vent speed to 10psi. 
    4. Steve says the vents typically take 4 of 300 cf cylinders from Airgas "Ultra Zero" AI  UZ300 that contains 0.1 PPM of THC

Everything looks good, so I'm monitoring the vent and swapping out cylinders.

At 12:08pm, the pressure was at 257 torr and I swapped out in a new cylinder.

Steve: Do not overpressurize the vacuum envelope! Stop around 720 Torr and let lab air do the rest. Our bellows are thin walled for seismic isolation.

Attachment 1: vent80wtiptilts.png  353 kB  Uploaded Fri Jul 13 14:45:31 2018  | Show | Show all
ELOG V3.1.3-